# Rethinking LLM-Based Recommendations: A Query Generation-Based, Training-Free Approach

### v1

@@ Abstract
Existing large language model (LLM)-based recommendation methods face several challenges, including inefficiency in handling large candidate pools, sensitivity to item order within prompts (“lost in the middle” phenomenon), poor scalability, and unrealistic evaluation with random negative sampling. To effectively address these issues, we propose a Query-to-Recommendation approach, which leverages LLMs to generate personalized queries for retrieving relevant items from the entire candidate pool, eliminating the need for candidate pre-selection. This method can be integrated into an ID-based recommendation system without additional training, enhances recommendation performance and diversity through LLM’s world knowledge, and works well even for minor item groups. Experiments on three datasets show up to 57% performance improvement, with an average gain of 31%, demonstrating competitive zero-shot performance and further gains when ensembled with existing models.
@@
A Query Generation-Based, Training-Free Approach
Donghee Han       Hwanjun Song       Mun Yong Yi KAIST, Republic of Korea {handonghee, songhwanjun, munyi}@kaist.ac.kr Corresponding author
@@ 1 Introduction
Recommendation systems play a crucial role in delivering personalized content and service across various domains, including e-commerce, social media, and streaming services. As the demand for more accurate, diverse and user-centric recommendations continues to increase, the integration of large language models (LLM) has emerged as a promising advancement Wu et al. (2024); Zhao et al. (2024). LLMs possess extensive knowledge of the world and exhibit remarkable abilities in understanding and generating text Yang et al. (2024), enabling new opportunities to improve recommendation systems beyond the limitations of traditional collaborative filtering (CF) and content-based methods. Recently, numerous studies have been proposed to leverage LLM in recommendation systems Ramos et al. (2024); Zhang et al. (2025); Lu et al. (2024).
However, existing LLM-based recommendation methods face significant challenges when applied to real-world environments. Most of the methods typically rely on incorporating candidate items directly into the prompt Hou et al. (2024); Kim et al. (2024); Bao et al. (2023); Liao et al. (2024), leading to inefficiencies, especially with large candidate pools that exceed 10K items. These approaches require a separate recommendation model for the selection of candidates from large-scale item pools. Additionally, the "lost in the middle" phenomenon occurs, where the position of items within the prompt affects the recommendation results Hou et al. (2024). Previous studies evaluated on naive random negative sampling do not accurately reflect real-world scenarios Hou et al. (2024); Bao et al. (2023); Kim et al. (2024); Bai et al. (2024). When an additional recommendation model is used for candidate selection, the selected items tend to exhibit higher similarity compared to random samples. It remains unclear whether LLMs can effectively rerank items in such cases.
@@
Another consideration when leveraging LLMs in recommendation systems is the cost and time required for training. Early study has proposed fine-tuning of LLMs on recommendation datasets to improve the recommendation ability Bao et al. (2023). More recent study used LLM as zero-shot ranker Hou et al. (2024); however, their performance was inferior compared to traditional training-based recommendation models. This result suggest that LLMs struggle to capture CF-signals, leading to the proposal of methods that enhance recommendation performance by learning CF-signals Kim et al. (2024); Dong et al. (2025). In these studies, user and item representations obtained from CF-based recommendation models were used for LLM training, and the training cost was reduced by fixing the LLM and updating only the weights of the additional projection layers. However, these methods still require gradient backpropagation through the LLM, which requires large GPU memory. Moreover, an additional training phase is needed to achieve alignment with the recommendation model.
Through our experiments, we observed that existing reranking-based methods struggle to perform effectively when using other recommendation models as candidate selectors. This indicates a critical need for a new approach that can operate efficiently within large candidate pools and seamlessly adapt to evolving LLMs without fine-tuning. The motivation behind this work lies in the unique strengths of LLMs. Their vast knowledge of the world enables them to provide flexible and context-sensitive recommendations in open-domain and unstructured scenarios Sun et al. (2024).
However, most studies primarily use LLMs as rerankers, which limits their potential by making them dependent on the performance of upstream candidate selectors. As a result, these methods can create bottlenecks due to the misalignment between the CF-based recommendation model and the differing perspectives of LLM. To address this issue, we propose a parallelized approach in which both the LLM and the recommendation model independently extract top-k items from the entire item pool.
@@
Fig.1 illustrates the difference between the traditional serialized approach and our proposed parallel approach. Our proposed parallel architecture maximizes the utilization of the LLM’s world knowledge and creates synergy by effectively merging the two complementary perspectives. We argue that LLMs can offer a complementary perspective to CF-based recommendation models. To fully leverage their potential, LLMs need direct access to the global item pool; however, incorporating the entire item pool into prompts is impractical. To address this, we introduce a query generation mechanism that enables the LLM to formulate feature-aware queries based on user’s history and key features of items.
To address these challenges, we propose Query-to-Recommendation (QueRec) framework, a novel recommendation method that leverages the complementary strengths of LLM and traditional recommendation models. Unlike existing serialized candidate selection LLM-reranking pipelines, our approach uses LLMs to generate personalized queries that retrieve relevant items directly from the global item pool. This zero-shot methodology eliminates the dependency on candidate pre-selection models and additional training, enabling flexible integration of new LLMs as they emerge. By combining CF signals from traditional recommendation models with world knowledge in LLMs, our method fosters a synergistic effect, enhancing both recommendation accuracy and diversity.
In our approach, the LLM generates feature-aware queries based on user history to incorporate its extensive background knowledge into the recommendation process, while the recommendation model contributes CF-based signals. These two components operate in a parallel architecture to enhance recommendation performance. Specifically, the LLM generates item queries from item information to create item embeddings and derives user queries from user history to produce user embeddings. The similarity between these embeddings is used to select top-k items. Finally, these items are reranked by integrating the similarity scores with the top-k items selected by the recommendation model, effectively combining both perspectives for improved recommendations.
@@
Our method not only improves the performance of the recommendation, but also promotes greater diversity and adaptability, effectively addressing the limitations of existing LLM-based recommendation systems. By generating feature-aware queries, it provides a more transparent rationale for item selection than approaches based on user reviews, offering direct insights into why specific items are recommended. The training-free nature of our framework allows for seamless incorporation of new items based on their key features without additional model updates, and it can also handle large, dynamically changing item pools. This advantage is particularly beneficial in real-world e-commerce scenarios, where large volumes of new items are continuously introduced and updated. Consequently, our design significantly reduces both infrastructure costs and maintenance overhead, making the system highly scalable for rapidly evolving recommendation environments.
@@
  • We propose a novel and practical Query-to-Recommendation approach that leverages LLMs to generate personalized queries, addressing the limitations of LLM-Reranking methods that rely on candidate sets.
  • Our method achieves competitive zero-shot performance comparable to state-of-the-art recommendation models, without requiring additional training.
  • We demonstrate that integrating our approach with existing methods improves performance, supported by experimental results.
@@ 2 Related Work
2.1 Sequential Recommendation Systems
@@ 2.2 LLM-based Recommendation Systems
With the rapid advancement of LLMs, there has been growing interest in incorporating them into recommendation systems Wu et al. (2024). Early LLM-based recommendation approaches fine-tuned pre-trained models such as T5 Raffel et al. (2020) to generate item recommendations by providing item or user IDs and outputting item IDs in a text format Geng et al. (2022); Li et al. (2023). However, this approaches had limitations in fully leveraging the rich knowledge embedded in LLMs. Recently, studies have proposed improving performance by distilling knowledge from larger LLMs Wang et al. (2024b) and incorporating CF-signals from user-item graphs to achieve better performance Wang et al. (2024a).
Another direction in LLM-based recommendation research explores natural language prompts instead of item and user IDs, allowing LLMs to leverage their extensive knowledge and zero-shot capabilities. Early studies fine-tuned LLMs using LoRA Bao et al. (2023), while later approaches froze LLM parameters and updated only additional components to incorporate CF-signals more efficiently Kim et al. (2024); Dong et al. (2025). Recent work introduced zero-shot LLM-based reranking Hou et al. (2024), but these methods remain limited by reliance on upstream candidate selectors, inefficiencies in large candidate pools, and the "lost in the middle" phenomenon. To address these challenges, we propose a method that better aligns with real-world recommendation settings, ensuring efficiency and scalability in large-scale item pools.
@@ 3 Method: QueRec Framework
3.1 Overview
We introduce Query-to-Recommendation (QueRec), a novel framework designed to leverage the complementary strengths of LLMs through large pool of items, while effectively connecting the diverse knowledge embedded in LLMs with the features of the elements. The proposed method leverages the world knowledge of LLMs to generate personalized queries, which can be effectively integrated with traditional recommendation models to improve the accuracy and diversity of recommendations. Fig.2 illustrates the overall process of our method. The overall process involves generating queries from item information to create item queries, leveraging user history to generate user queries, retrieving top-k items based on similarity scores between user queries and item queries, and finally applying a reranking mechanism that fuses insights from both the LLM and the recommendation model. This architecture enables robust, diverse, and personalized recommendations without the need for extensive training of LLMs, and direct search is available in a large item pool, making it scalable for real-world applications.
3.2 Item Query Generation
To effectively leverage the world knowledge of LLMs to highlight key features of items and align them with user preferences, we employ LLMs to generate queries that reflect the distinctive features of each item. This process involves constructing prompts using item metadata (e.g. title, category, description) and user reviews to provide contextual information to the LLM. Based on these prompts, the LLM generates 10 diverse queries that capture different aspects of the item, leveraging its extensive knowledge. Finally, the generated queries are combined with the original item metadata to create a comprehensive item representation.
@@ 3.3 User Query Generation
The user representation is generated by incorporating information from the user’s historical interactions and preferences. This process begins with the construction of prompts that combine user history (including item titles and metadata), user reviews, and previously generated item queries to provide rich contextual information. Based on these prompts, the LLM generates 10 personalized queries that reflect the user’s interaction history. The generated queries are then refined by incorporating the most recently interacted item information, ensuring that recent user preferences are effectively captured.
$q_{u}=\text{LLM}(p_{u},h_{u},r_{u},q_{l}),\quad d_{u}=m_{l}\oplus q_{u}$ (2)
Equation (2) describes the process of constructing the textual user representation. Given a user $u$, we construct a query $q_{u}$ using the LLM with inputs: user query generation prompt $p_{u}$, user history $h_{u}$ (which includes item titles and metadata), user reviews $r_{u}$, and the most recent item query $q_{l}$. The final user representation $d_{u}$ is obtained by concatenating the user history with the generated query and incorporating the last interacted item metadata $m_{l}$ to better reflect temporal user preferences. The prompt format and examples are described in the Appendix C .
@@ 3.4 Similarity-based Retrieval
With the user and item textual representations obtained from the previous steps, we perform similarity-based retrieval:
$v_{u}$ $=\text{Enc}(d_{u}),\quad v_{i}=\text{Enc}(d_{i})$ (3)
@@
^$\hat{\mathcal{I}}_{u}=\underset{i\in\mathcal{I}}{\text{arg max }s_{u,i}},\quad\text{where }|\hat{\mathcal{I}}_{u}|=k$ (4)
Equation (3) defines the embedding extraction via the pre-trained text encoder Enc and similarity computation between the user embedding $v_{u}$ and item embeddings $v_{i}$. We compute cosine similarity $s_{u,i}$ between the embeddings to measure their relevance. Based on these scores, we select the top-k items ^$\hat{\mathcal{I}}_{u}$ for the user query, as shown in Equation (4), which are the most relevant items to the user’s preferences.
@@ 3.5 Divergent Perspective Reranking
To effectively integrate LLM-based semantic insights and CF-based collaborative signals, we propose a Divergent Perspective Reranking method. Initially, similarity scores from both models (LLM and CF) are normalized into the [0,1] range via min-max normalization, ensuring fair comparison across methods:
@@
Subsequently, we calculate the final score using a convex combination (CC) of the normalized scores:
$s_{u,i}^{*}=\lambda\tilde{s}_{u,i}^{LLM}+(1-\lambda)\tilde{s}_{u,i}^{CF}.$ (6)
@@
$\lambda_{\text{init}}=\frac{H_{10}^{LLM}}{H_{10}^{LLM}+H_{10}^{CF}}.$ (7)
Further analysis revealed that the intersection ratio ($\omega$) between hit items from QueRec and CF models varies significantly across datasets and models. In the case of low intersection ratio with imbalanced weights (e.g., 0.3) causes interference between the two recommended item lists, resulting in performance decreasing. To address this, we adjust $\lambda$ by incorporating the intersection ratio:
$\lambda=\omega\cdot\lambda_{\text{init}}+(1-\omega)\cdot 0.5.$ (8)
When overlap is minimal ($\omega\approx 0$), $\lambda$ naturally converges to 0.5, reflecting uncertainty and ensuring a balanced influence from both sources. This adjustment enhances recommendation quality by effectively merging diverse perspectives from LLM and CF knowledge.

### v2

# Rethinking LLM-Based Recommendations: A Personalized Query-Driven Parallel Integration

@@ Abstract
Recent studies have explored integrating large language models (LLMs) into recommendation systems but face several challenges, including training-induced bias and bottlenecks from serialized architecture. To effectively address these issues, we propose a Query-to-Recommendation, a parallel recommendation framework that decouples LLMs from candidate pre-selection and instead enables direct retrieval over the entire item pool. Our framework connects LLMs and recommendation models in a parallel manner, allowing each component to independently utilize its strengths without interfering with the other. In this framework, LLMs are utilized to generate feature-enriched item descriptions and personalized user queries, allowing for capturing diverse preferences and enabling rich semantic matching in a zero-shot manner. To effectively combine the complementary strengths of LLM and collaborative signals, we introduce an adaptive reranking strategy. Extensive experiments demonstrate an improvement in performance up to 57%, while also improving the novelty and diversity of recommendations.
@@
A Personalized Query-Driven Parallel Integration
Donghee Han       Hwanjun Song       Mun Yong Yi KAIST, Republic of Korea {handonghee, songhwanjun, munyi}@kaist.ac.kr Corresponding authors
@@ 1 Introduction
Recommendation systems play a crucial role in delivering personalized content and service across various domains. As the demand for more accurate and diverse recommendations continues to increase, the integration of large language models (LLM) has emerged as a promising advancement Wu et al. (2024); Zhao et al. (2024). LLMs possess extensive knowledge and exhibit remarkable abilities in understanding and generating text Yang et al. (2024), enabling new opportunities to improve recommendation systems beyond the traditional collaborative filtering (CF) and content-based methods. Recently, numerous studies have been proposed to leverage LLM in recommendation systems Ramos et al. (2024); Zhang et al. (2025a); Lu et al. (2024).
To harness the capabilities of LLMs for recommendation, two major research paradigms have emerged. The first line of works utilize LLMs as generative predictors, typically fine-tuning them on next-item prediction tasks. These methods generate textual representations (e.g., item titles) based on user histories and use them as queries for retrieving candidate items Bao et al. (2025); Li et al. (2023b). While this approach leverages LLMs’ generation strength, it requires fine-tuning, which introduces bias to train dataset and reduces diversity of recommendation. Additionally, relying solely on item titles can limit the expressive capacity of LLMs, limiting their ability to align nuanced user preferences with rich item characteristics.
The second line of research focuses on reranking candidate items directly within the LLM prompt. In this setting, a separate candidate retrieval model selects a subset of items, which are then presented to the LLM for scoring or reranking Hou et al. (2024); Kim et al. (2024); Bao et al. (2023). These approaches typically leverage CF-based recommendation models such as SASRec Kang and McAuley (2018) for candidate selection Yang et al. (2023), and various techniques have been proposed to align LLMs with CF signals Kim et al. (2024); Dong et al. (2025).
However, these approaches inherently depend on the performance of the candidate selector, and they limit the ability of LLMs to fully leverage their distinctive capabilities across the entire item pool. The information handled by the CF-based model and the LLM is fundamentally different, and the current serialized architecture can limit performance due to the misalignment between these two components. This misalignment also becomes a bottleneck that hinders the utilization of the diverse knowledge embedded in LLMs for recommendation.
@@
In this study, we focus on addressing two key challenges in LLM-based recommendation: the bias introduced by fine-tuning LLMs on recommendation datasets, and the bottleneck caused by the serialized architecture that utilizes CF-based models as candidate selectors. Motivated by the fact that CF-based models and LLMs handle fundamentally different types of information, we propose a parallel framework that integrates both components without entangling their processes. Our approach enables the generation of enriched, feature-aware queries that allow LLMs to effectively leverage their broad and diverse knowledge for recommendation.
@@
Fig.1 illustrates the difference between the traditional serialized approach and our proposed parallel approach. Our proposed parallel architecture maximizes the utilization of the LLM’s world knowledge and creates synergy by effectively merging the two complementary perspectives. We argue that LLMs can offer a complementary perspective to CF-based recommendation models. To fully leverage their potential, LLMs need direct access to the global item pool; however, existing approaches that fine-tune LLMs to generate item titles hinder the utilization of the diverse knowledge of LLMs and introduce biases from the training dataset..
In our framework, LLMs and CF-based models independently retrieve top-k items from the entire candidate pool. By decoupling LLMs from pre-selected candidates, our method maximizes LLMs’ ability and creates synergy between complementary perspectives. To utilize LLMs’ rich world knowledge, we introduce item description expansion and personalized query generation where the LLM generates feature-enriched item description and personalized user queries grounded in user histories and item characteristics. Furthermore, we design an adaptive reranking strategy that aggregates the outputs of LLMs and CF-based models while maintaining their independency. Unlike existing approaches that jointly train multiple models, our method dynamically balances contributions.
@@
Our method not only improves the performance of the recommendation, but also promotes greater diversity and adaptability, effectively addressing the limitations of existing LLM-based recommendation systems. The training-free nature of our framework allows for seamless incorporation of new items based on their key features without additional model updates, and it can also handle large, dynamically changing item pools. This advantage is particularly beneficial in real-world e-commerce scenarios, where large volumes of new items are continuously introduced and updated. Consequently, our design significantly reduces both infrastructure costs and maintenance overhead, making the system highly scalable for rapidly evolving recommendation environments.
Figure 2: Overview of QueRec.
@@
  • We introduce a parallel architecture for recommendation that combines LLM-based retrieval and CF-based models, maximizing the complementary strengths of each component.
  • Our method outperforms state-of-the-art methods and achieved up to 57% improvement in performance over traditional CF-based models, without requiring additional LLM fine-tuning.
  • Through our analysis, we demonstrate that our approach effectively addresses bias and diversity issues in recommendations.
@@ 2 Related Work
2.1 Traditional Recommendation Systems
@@ 2.2 LLM-based Recommendation Systems
With the rapid advancement of LLMs, there has been a growing interest in incorporating them into recommendation systems Wu et al. (2024). Early LLM-based recommendation approaches fine-tuned pre-trained models such as T5 Raffel et al. (2020) to generate item recommendations by providing item or user IDs and output item IDs in a text format Geng et al. (2022); Li et al. (2023c); Wang et al. (2024b, a). However, these approaches had limitations in fully leveraging the rich knowledge embedded in LLMs.
Another direction in LLM-based recommendation research explores natural language prompts instead of item and user IDs, allowing LLMs to leverage their extensive knowledge and zero-shot capabilities. Early studies fine-tuned LLMs Bao et al. (2023), while later approaches fixed LLM parameters and update only additional components to incorporate CF-signals Kim et al. (2024); Dong et al. (2025). Recent work introduced zero-shot LLM-based reranking Hou et al. (2024), but these methods remain limited by the reliance on upstream candidate selectors and positional bias.
Recently, other studies proposed fine-tuning LLMs to generate item titles and using the generated titles as search queries for recommendationLi et al. (2023b); Bao et al. (2025). However, these approaches have limitations in leveraging the diverse knowledge inherent in LLMs. To address these challenges, we propose a method that better aligns with real-world recommendation settings, ensuring ease of deployment and scalability in large-scale item pools.
@@ 3 Method: QueRec Framework
In this section, we introduce QueRec(Query-to-Recommendation), a novel framework designed to leverage the LLMs through large item pool and to operate in parallel with CF-based models, enabling a dual-perspective that integrates both collaborative signals and language-based knowledge. Fig.2 illustrates the overall process of our method. The proposed method leverages LLMs to generate feature-enriched item descriptions and personalized user queries, retrieving top-k items based on similarity scores between user queries and item descriptions. These LLM-based results are then combined with the top-k items from a CF-based model, effectively fusing insights from both CF-based models and LLMs.
3.1 Item Database Construction
To effectively leverage the world knowledge of LLMs to highlight key features of items and align them with user preferences, we employ LLMs to generate feature-enriched item descriptions that reflect the distinctive features of each item. Fig.2(a) illustrates the process of leveraging an LLM to expand item descriptions, which are then encoded into embeddings to construct an item vector database.
This process involves constructing prompts using item metadata (e.g. title, category, description) and user reviews to provide contextual information to the LLM. We design the prompt to expand item descriptions into a set of diverse queries, aiming to better reflect users’ varied preferences. Given these prompts, the LLM generates ten distinct queries, each capturing different aspects of the item by leveraging its extensive knowledge. Finally, the generated queries are combined with the original item metadata to create a comprehensive item representation, which is then converted into an item embedding using a retrieval encoder and cached for search.
@@ 3.1 Item Database Construction
$\mathcal{V}=\{\text{Enc}(d_{i})\mid d_{i}\in\mathcal{D}\}$ (2)
Continuing from the previous step, we encode all enriched item descriptions $d_{i}\in\mathcal{D}$ using a retrieval encoder and cache the resulting embeddings $\mathcal{V}$ for efficient search during recommendation, as shown in Equation (2).
@@ 3.2 User Query Generation
To represent the user, we incorporate information from their historical interactions and preferences. This process begins with the construction of prompts that combine user history (including item titles and metadata), user reviews, and previously generated item descriptions to provide rich contextual information. Based on these prompts, the LLM generates ten personalized queries that reflect the user’s context and preferences. To emphasize the user’s most recent preferences, we append the enriched description of the most recently purchased item to the prompt.
$q_{u}=\text{LLM}(p_{u},h_{u},r_{u},d_{l}),\quad d_{u}=m_{l}\oplus q_{u}$ (3)
Equation (3) describes the process of constructing the textual user representation. Given a user $u$, we construct a query $q_{u}$ using the LLM with inputs: user query generation prompt $p_{u}$, user history $h_{u}$ (which includes item titles and metadata), user reviews $r_{u}$, and the most recent item description $d_{l}$. The final user representation $d_{u}$ is obtained by concatenating the user history with the generated query and incorporating the last interacted item metadata $m_{l}$ to better reflect temporal user preferences.
An important aspect of this step is that, instead of using each generated query independently, we aggregate them into a single unified user representation. Similar to the item description expansion, we prompt the LLM to generate multiple personalized queries in order to capture diverse user perspectives and preferences. These queries are not treated separately but are concatenated into a single user representation, which is then used to construct a comprehensive user embedding. The prompt format and generated query examples are described in the Appendix G.
@@ 3.3 Similarity-based Retrieval
With the user textual representations and item database obtained from the previous steps, we perform similarity-based retrieval for personalized recommendation :
$v_{u}$ $=\text{Enc}(d_{u}),\quad v_{i}\in\mathcal{V}$ (4)
@@
$\hat{\mathcal{I}}_{u}=\underset{i\in\mathcal{I}}{\text{arg max }s_{u,i}},\quad\text{where }|\hat{\mathcal{I}}_{u}|=k$ (5)
Equation (4) defines the embedding extraction via the pre-trained text encoder Enc and similarity computation between the user embedding $v_{u}$ and item embeddings in item database $v_{i}\in\mathcal{V}$. We compute cosine similarity $s_{u,i}$ between the embeddings to measure their relevance. Based on these scores, we select the top-k items $\hat{\mathcal{I}}_{u}$ for the user query, as shown in Equation (5), which are the most relevant items to the user’s preferences.
Our semantic similarity-based retrieval utilizing personalized queries offers richer information compared to traditional grounding methods that rely solely on item titles. It is also more adaptable to updates in the item pool, which occur frequently in real-world recommendation scenarios.
@@ 3.4 Divergent Perspective Reranking
To effectively integrate LLM-based semantic insights and CF-based collaborative signals, we propose a divergent perspective reranking method. Initially, similarity scores from each model (LLM and CF) are normalized, ensuring fair influence from each method:
@@
In Equation (6), $s_{u,i}^{X}$ and $\tilde{s}_{u,i}^{X}$ denote the original score and final normalized score, between user $u$ and item $i$ computed by model $X$ (LLM-based and CF-based). For each user, we apply min-max normalization over the score set $s_{u}^{X}=\{s_{u,i}^{X}\mid i\in\mathcal{I}_{u}\}$, where $\mathcal{I}_{u}$ represents the entire item pool. This normalization maps all user-item relevance scores to the range $[0,1]$, ensuring that the scores from different models are comparable on the same scale. The process preserves the internal ranking structure of each model while aligning their score magnitudes for reranking.
Subsequently, we calculate the final reranking score using a convex combination (CC) of the normalized scores:
$\vskip-2.84526pts_{u,i}^{*}=\lambda\tilde{s}_{u,i}^{LLM}+(1-\lambda)\tilde{s}_{u,i}^{CF}.$ (7)
@@
$\vskip-2.84526pt\lambda_{\text{init}}=\frac{H_{10}^{LLM}}{H_{10}^{LLM}+H_{10}^{CF}}.$ (8)
This method of computing the initial $\lambda$ is intended to assign a larger weight to the model with higher performance. However, it may be less effective in cases where the two models perform well on entirely different user groups.
Further analysis revealed that the intersection ratio between hit items from QueRec and CF models varies significantly across datasets and models. In the case of a low intersection ratio with imbalanced weights (e.g., 0.3), interference between the two recommended item lists causes performance to decrease. To address this, we adjust $\lambda$ by incorporating the intersection ratio:
$\vskip-5.69054pt\lambda=\omega\cdot\lambda_{\text{init}}+(1-\omega)\cdot 0.5.$ (9)
Here, $\omega$ is computed as:
$\vskip-2.84526pt\omega=\frac{|\mathcal{H}_{10}^{\text{LLM}}\cap\mathcal{H}_{10}^{\text{CF}}|}{|\mathcal{H}_{10}^{\text{LLM}}\cup\mathcal{H}_{10}^{\text{CF}}|},$ (10)
where $\mathcal{H}_{10}^{\text{LLM}}$ and $\mathcal{H}_{10}^{\text{CF}}$ denote the sets of user sequences whose top-10 lists generated by the LLM-based and CF-based models, respectively, contain the target item.
When the intersection ratio is minimized ($\omega\approx 0$), $\lambda$ converges to 0.5, reflecting the orthogonality between the two models, and ensuring a balanced influence from both models. This adjustment improves the quality of the recommendation by effectively merging diverse perspectives from the LLM and CF-based method. The analysis related to this approach is explained in Appendix D.